# Reasoning Grasping via Multimodal Large Language Model

### v1

@@ V-B1 Part Segmentation
The annotation of object parts was a critical step in our dataset creation. We manually segmented the part point clouds of each object using a specially developed annotation tool. The segmentation process was guided by an understanding of how humans typically interact with and use various parts of an object. For instance, a knife was divided into its blade and handle, reflecting the distinct functional roles of each part.
@@ V-B2 Grasping Pose Annotation
Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses of each object in the scene.
Pose Assignment. Pose assignment involves aligning the grasping poses with the nearest part of the object. For each object’s segmented parts, we determine the grasping center of each point. These grasping poses are then assigned to the nearest part, ensuring an accurate association with the specific object part they correspond to.
Pose Projection. Pose projection entails transforming the grasping poses of parts into both 3D and 2D representations. This is achieved using camera matrices and object transformations. In the case of 3D, the original grasping poses are preserved in the spatial domain. For 2D projection, the 3D points of each segmented part are projected onto 2D images. This results in 2D masks that depict the spatial distribution of each part on the image plane. Grasping poses within these masks are then mapped to the corresponding parts in 2D.
These two processes of pose assignment and projection form a comprehensive approach to annotating our dataset. This approach ensures that the grasping pose data is not only detailed and extensive but also accurately aligned with the physical and functional characteristics of each object.
@@ V-C Dataset Statistics
Our dataset significantly enhances the GraspNet-1B foundation by annotating object parts and incorporating reasoning instructions for robotic grasping tasks and detailed object parts annotations.
In ensuring data quality, low-quality grasping poses were eliminated based on confidence evaluations and objects lacking semantic information were removed. In total, our enhanced dataset comprises an extensive collection of 1,730 instruction pairs, 64 objects, 109 segmented parts, and around 100 million associated grasping poses.
Table I shows the statistics comparison of the previous grasping and visual reasoing datasets with ours. Our dataset distinguishes itself from existing visual reasoning works by integrating actionable robotic manipulation data, a feature not commonly found in similar datasets. Moreover, unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information. This unique combination of features positions our dataset as a critical resource for developing more intelligent and adaptable robotic systems, bridging the gap between visual reasoning and practical robotic tasks.
@@ VI-A Implementation Details
For the multi-modal LLM in the proposed model, we utilize the pre-trained LLaVA-7B-v0, which is derived from the large language model LLaMA-7B. As for the LoRA fine-tuning, we choose a rank $r=64$. During the training, we use the batch size of $8$, and the learning rate is set to $45e-$ 4, utilizing a cosine annealing schedule for adaptive learning rate adjustment. For the relative weight parameters in the loss function 2, we set $\lambda_{t}=1$ and $\lambda_{g}=1$ for the initial configuration, and starting from the third epoch, the parameter $\lambda_{t}$ is adjusted to 00.
@@ VI-B2 LLaVa $\rightarrow$ GR-ConvNet Baseline
This is a modular baseline that integrates the latest LLaVA-1.6-34b [26] with a pre-trained grasping detection model GR-ConvNet. As LLaVA cannot directly output the grasping pose, this baseline operates in two stages to obtain the grasping pose: first, LLaVA takes an image and an instruction as the input and is prompted to output a bounding box of the target object in the image. Then GR-ConvNet detects the optimal grasping pose within this specified bounding box. Note that LLaVA-1.6-34b has a much higher performance than LLaVa-v0-7b [25], which is the base model of our proposed reasoning grasping model.
@@ VI-C1 Text Generation
Models
R@1
R@2
R@3
R@1
R@2
R@3
R@1
R@2
R@3
R@1
R@2
R@3
LLaVA $\rightarrow$ GR-ConvNet
26.80%
39.17%
44.66%
18.34%
23.76%
31.29%
15.38%
23.52%
37.93%
0.92%
3.84%
8.53%
CLIP $+$ GR-ConvNet
50.40%
62.88%
66.38%
44.35%
57.15%
61.09%
43.78%
55.71%
59.61%
28.88%
42.34%
47.37%
@@
LLaVA $\rightarrow$ GR-ConvNet (with depth)
31.88%
40.72%
45.95%
20.39%
25.77%
34.84%
20.11%
29.30%
36.63%
1.36%
2.77%
11.95%
CLIP $+$ GR-ConvNet (with depth)
54.11%
66.50%
69.57%
45.89%
58.11%
62.50%
45.53%
59.51%
62.27%
33.06%
48.55%
52.25%
@@ VI-C3 Visual Reasoning Ability
The image-question-answer sets used in this evaluation are randomly chosen from the LLaVa-Instruct-150k dataset [25]. According to the judge of GPT-4, the original LLaVA-7B-v0 achieves $8.25/10$ when taking ground-truth textual descriptions as references. On the other hand, our model achieves $7.43/10$ under the same GPT-4 judge. This shows that our fine-tuned model still preserves a good portion of the reasoning ability of the original LLaVA.
@@ VI-D Real World Experiments
Overall our proposed model can generate accurate special tokens and target names 39 out of 40 trials. This again shows the excellent reasoning capabilities of our model to interpret the various instructions. The failure case involved the model incorrectly identifying a screwdriver instead of scissors in response to a prompt requesting a tool to open a package box. In terms of grasping performance, we assessed top-3 output grasp poses and manually selected the most feasible one for execution on the robot arm. The CLIP $+$ GR-ConvNet baseline and our reasoning grasping model produced 15/40 and 25/40 accurate grasping poses, respectively, with the success rates for lifting objects are 9/40 and 16/40, respectively. The experiment results demonstrate that our model outperforms the baseline in four distinct scenarios, showing its superior ability in reasoning grasping tasks.

### v2

@@ V-B1 Part Segmentation
We manually segmented the part point clouds of each object using a specially developed annotation tool. The segmentation process was guided by an understanding of how humans typically interact with and use various parts of an object. For instance, a knife was divided into its blade and handle, reflecting the distinct functional roles of each part.
@@ V-B2 Grasping Pose Annotation
Our dataset’s annotation process includes two key components: pose assignment and pose projection, for both 3D and 2D grasping poses, as shown in Fig. 4.
Pose assignment involves aligning the grasping poses with the nearest part of the object. For each object’s segmented parts, we determine the grasping center of each point. These grasping poses are then assigned to the nearest part, ensuring an accurate association with the specific object part they correspond to.
Pose projection transforms the grasping poses of parts into both 3D and 2D representations. This is achieved using camera matrices and object transformations. In the case of 3D, the original grasping poses are preserved in the spatial domain. For 2D projection, the 3D points of each segmented part are projected onto 2D images. This results in 2D masks that depict the spatial distribution of each part on the image plane. Grasping poses within these masks are then mapped to the corresponding parts in 2D.
@@ V-C Dataset Statistics
In ensuring data quality, low-quality grasping poses were eliminated based on confidence evaluations and objects lacking semantic information were removed. In total, our enhanced dataset comprises an extensive collection of 1,730 instruction pairs, 64 objects, 109 segmented parts, and around 100 million associated grasping poses. Table I shows the statistics comparison of the previous grasping and visual reasoning datasets with ours. Unlike other grasping datasets, we offer detailed reasoning instructions and part-specific grasping information.
@@ VI-A Implementation Details
For the multi-modal LLM in the proposed model, we utilize the pre-trained LLaVA-7B-v0, which is derived from the large language model LLaMA-7B. As for the LoRA fine-tuning, we choose a rank $r=64$. During the training, we use the batch size of $8$, and the learning rate is set to $45e-$ 4, utilizing a cosine annealing schedule for adaptive learning rate adjustment. For the relative weight parameters in the loss function 1, we set $\lambda_{t}=1$ and $\lambda_{g}=1$ for the initial configuration, and starting from the third epoch, the parameter $\lambda_{t}$ is adjusted to 00. For the train-test-split, we use $90\%$ of the first 100 scenes in GraspNet-1 billion [13] for training and $10\%$ of the first 100 scenes for testing. We sampled 50k from the original LLaVA Instruct 150K dataset [24] to mix with 100k grasping data to maintain the visual reasoning capabilities.
@@ VI-B2 LLaVa $\rightarrow$ GR-ConvNet Baseline
This is a modular baseline that integrates the latest LLaVA-1.6-34b [25] with a pre-trained grasping detection model GR-ConvNet. As LLaVA cannot directly output the grasping pose, this baseline operates in two stages to obtain the grasping pose: first, LLaVA takes an image and an instruction as the input and is prompted to output a bounding box of the target object in the image. Then GR-ConvNet detects the optimal grasping pose within this specified bounding box. Note that LLaVA-1.6-34b has a much higher performance than LLaVa-v0-7b [24], which is the base model of our proposed reasoning grasping model. This baseline mirrors a similar process to LAN-grasp.
@@ VI-C1 Text Generation
Models R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3 R@1 R@2 R@3
LLaVA $\rightarrow$ GR-ConvNet 26.80% 39.17% 44.66% 18.34% 23.76% 31.29% 15.38% 23.52% 37.93% 0.92% 3.84% 8.53%
CLIP $+$ GR-ConvNet 50.40% 62.88% 66.38% 44.35% 57.15% 61.09% 43.78% 55.71% 59.61% 28.88% 42.34% 47.37%
@@
LLaVA $\rightarrow$ GR-ConvNet (with depth) 31.88% 40.72% 45.95% 20.39% 25.77% 34.84% 20.11% 29.30% 36.63% 1.36% 2.77% 11.95%
CLIP $+$ GR-ConvNet (with depth) 54.11% 66.50% 69.57% 45.89% 58.11% 62.50% 45.53% 59.51% 62.27% 33.06% 48.55% 52.25%
@@ VI-C3 Visual Reasoning Ability
The image-question-answer sets used in this evaluation are randomly chosen from the LLaVa-Instruct-150k dataset [24]. According to the judge of GPT-4, the original LLaVA-7B-v0 achieves $8.25/10$ when taking ground-truth textual descriptions as references. On the other hand, our model achieves $7.43/10$ under the same GPT-4 judge. This shows that our fine-tuned model still preserves a good portion of the reasoning ability of the original LLaVA. It also shows that the multi-round conversation ability is preserved, as the image-question-answer sets include sequences of multi-round conversations.
@@ VI-D Real World Experiments
Overall our proposed model can generate accurate special tokens and target names 39 out of 40 trials. This again shows the excellent reasoning capabilities of our model to interpret the various instructions. The failure case involved the model incorrectly identifying a screwdriver instead of scissors in response to a prompt requesting a tool to open a package box. We classified it as a failure solely because it did not match the specific pairing in our dataset, which serves as the basis for our evaluation metrics. This example underscores that even when the model’s response is technically incorrect, it is still reasonable. In terms of grasping performance, we assessed top-3 output grasp poses and manually selected the most feasible one for execution on the robot arm. We conducted 10 trials for each scenario, totaling 80 trials, to compare reasoning grasping with CLIP $+$ GR-ConvNet baseline. The CLIP $+$ GR-ConvNet baseline and our reasoning grasping model produced 15/40 and 25/40 accurate grasping poses, respectively, with the success rates for lifting objects are 9/40 and 16/40, respectively. The main reasons for accurate poses but fail to lift were due to the object slipping from the gripper or the object being too heavy. The experiment results demonstrate that our model outperforms the baseline in four distinct scenarios, showing its superior ability in reasoning grasping tasks.